OpenStudy (anonymous):

2 + 4(x - 3) + 2 = 4 + 5y in standard form? Medal and Fan :D

OpenStudy (anonymous):

\[y=\frac{ 4 }{ 5 }x-\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

How did you get that??

OpenStudy (haleyelizabeth2017):

I thought standard form was Ax+By=c

OpenStudy (anonymous):

Oh so move the x's to one side then the y's then the answer?

OpenStudy (haleyelizabeth2017):

I can walk you through how to get it if you'd like?

OpenStudy (anonymous):

yeah, sorry i got it wrong

OpenStudy (anonymous):

Please do @haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

Okay, so what we want to do first is distribute the 4 to the parenthesis (x-3). Which gives us...?

OpenStudy (anonymous):

2 + 4*x + 12 + 2 = 4 + 5y ?

OpenStudy (haleyelizabeth2017):

It would actually be -12 :) So 2+4x-12+2=4+5y

OpenStudy (haleyelizabeth2017):

Then, we want to combine like terms. So 2+2-12 is what?

OpenStudy (anonymous):

8

OpenStudy (haleyelizabeth2017):

-8 actually

OpenStudy (haleyelizabeth2017):

So... -8+4x=4+5y, then we add 8 to both sides to get 4x=12+5y Then we want to subtract 5y from both sides to get it in standard form. This gives us 4x-5y=12 as our final answer :)

OpenStudy (anonymous):

Thanks